# Structured Representation

### v1

@@ 4 Discussion
Figure 4: Illustration of partition block Structures applied to feasible relations in a navigation environment (Inset). Axes are color-coded for clarity: the first row and column adjacent to the axes indicate the type, as defined by the accompanying colormap, while the second row and column denote the entity color. Left: Jordan block partitions, where each block is distinctly highlighted. The source and sink nodes of each block are at the top and bottom, respectively. Right: Maximal antichain nodes are marked with red circles while their corresponding zeta blocks over the reflexive transitive closure are highlighted in lime overlaying closure.
We uncover partition block structures, maximal antichains and the corresponding zeta blocks over the reflexive transitive closure, illustrated in Figure 4, by application to a feasible relation matrix in a navigation domain, the MiniGrid obstructed maze environment Chevalier-Boisvert et al. (2018). The environment comprises nine rooms and objects such as balls, boxes, keys, and doors; some doors are locked requiring matching keys. The feasible relation matrix is precomputed, encoding feasible binary relations between entity attributes. The considered relations include: ‘in_room’ - entity rooms; ‘is_adj’ and ‘can_be_moved’ - movable entities balls, keys, and boxes which are either already adjacent or can be moved within the same room; ‘can_contain’ - a box can contain a key; and ‘open_door’ - a key can be used to open a door if it matches in color, or the agent can open an unlocked door.

### v2

@@ 4 Discussion
Figure 4: Visualization of partition block structures uncovered in application to feasible relations in a navigation environment (Inset). Axes are color-coded for clarity: the first row and column adjacent to the axes indicate the type, as defined by the accompanying colormap, while the second row and column denote the entity color. Left: Jordan block partitions, where each block is distinctly highlighted. The source and sink nodes of each block are at the top and bottom, respectively. Right: Maximal antichain nodes are marked with red circles while their corresponding zeta blocks over the reflexive transitive closure are highlighted in lime overlaying closure.
We uncover partition block structures, maximal antichains and the corresponding zeta blocks over the reflexive transitive closure, shown in Figure 4, by application to a feasible relation matrix in a navigation domain, the MiniGrid obstructed maze environment Chevalier-Boisvert et al. (2018). The environment comprises nine rooms and objects such as balls, boxes, keys, and doors; some doors are locked requiring matching keys. The feasible relation matrix is precomputed, by encoding feasible binary relations between entity attributes. The considered relations are: ‘in_room’ - entity rooms; ‘is_adj’ and ‘can_be_moved’ - movable entities balls, keys, and boxes which are either already adjacent or can be moved within the same room; ‘can_contain’ - a box can contain a key; and ‘open_door’ - a key can be used to open a door if it matches in color, or the agent can open an unlocked door. The feasible relation matrix is an attributes-attributes binary relation matrix as described above. For the relation feasibility matrix, the Jordan blocks structure is revealed using SymPy. The relational reflexive transitive closure is computed using the Floyd-Warshall algorithm, while the maximal antichains are obtained by checking the reachability and then used to obtain the zeta block structures, see Figure 4.
@@
This work focuses on developing foundations of invariant knowledge partitions that support knowledge learning and representations; low-level navigation lies beyond the scope of this article.